# Controllable Logical Hypothesis Generation for Abductive Reasoning in Knowledge Graphs

### v1

@@ Abstract
Abductive reasoning in knowledge graphs aims to generate plausible logical hypotheses from observed entities, with broad applications in areas such as clinical diagnosis and scientific discovery. However, due to a lack of controllability, a single observation may yield numerous plausible but redundant or irrelevant hypotheses on large-scale knowledge graphs. To address this limitation, we introduce the task of controllable hypothesis generation to improve the practical utility of abductive reasoning. This task faces two key challenges when controlling for generating long and complex logical hypotheses: hypothesis space collapse and hypothesis oversensitivity. To address these challenges, we propose CtrlHGen, a Controllable logcial Hypothesis Generation framework for abductive reasoning over knowledge graphs, trained in a two-stage paradigm including supervised learning and subsequent reinforcement learning. To mitigate hypothesis space collapse, we design a dataset augmentation strategy based on sub-logical decomposition, enabling the model to learn complex logical structures by leveraging semantic patterns in simpler components. To address hypothesis oversensitivity, we incorporate smoothed semantic rewards including Dice and Overlap scores, and introduce a condition-adherence reward to guide the generation toward user-specified control constraints. Extensive experiments on three benchmark datasets demonstrate that our model not only better adheres to control conditions but also achieves superior semantic similarity performance compared to baselines.
@@ 1 Introduction
Abduction is widely recognized as one of the three major types of reasoning in philosophy abduction . Specifically, abductive reasoning abduction is a form of logical inference that seeks the best or most plausible hypothesis to explain an observed phenomenon and it plays a vital role across various fields abducreview . For example, it serves as a critical tool for hypothesizing causal links between symptoms and underlying pathologies in clinical diagnosis medicaldianosis ; medical2 . Similarly, abductive methods localize system faults by interpreting anomalous signal patterns in anomaly detection attack ; attack2 . Its power also extends to scientific discovery science1 ; science2 ; science3 ; science4 , including the deduction of unknown celestial bodies from gravitational perturbations in orbital trajectories sciencediscovery .
@@
On the other hand, effective abductive reasoning requires high-quality interconnected information as premises. Knowledge graphs, whether general or domain-specific, provide an ideal structure for these premises, supporting robust abductive reasoning processes. Abductive reasoning in knowledge graphs aims to generate complex logical hypotheses that best explain a given set of observed entities, while maintaining consistency with the existing knowledge base. This process leverages the rich domain knowledge, thereby improving both the precision and reliability of the inference. AbductiveKGR akgr is the first work to realize abductive reasoning over knowledge graphs, training a hypothesis generation model using a supervised learning–reinforcement learning paradigm.
However, as knowledge graphs often contain millions of facts, an observation can lead to numerous diverse yet plausible hypotheses, most of which may be irrelevant to the specific concerns. To address this challenge, we are motivated to introduce controlling mechanisms in the hypothesis generation process that focus on two critical aspects:
@@
Controlling structural complexity adjusts the level of granularity. We focus on complexity control to address varying information needs across different reasoning scenarios and align with users’ cognitive preferences for adjustable information density. In Fig. 1(b), for an observation composed of three NBA players, increasing the complexity of the hypothesis structure enables the model to capture richer shared experiences or achievements among them. By adjusting the structural complexity, users can flexibly decide how much information they want to include in the generated hypotheses. Unfortunately, prior work akgr has largely overlooked controllable generation, resulting in hypotheses that are redundant or lack meaningful relevance.
@@
Motivated by these, we introduce the task of controllable abductive reasoning, aiming at controllable generation of hypothesis, which leads to better leverage the practical value of abductive reasoning in knowledge graphs. However, when implementing semantic and structural controls on complex logical hypotheses, we face two critical challenges: (i) Hypothesis Space Collapse: As illustrated in Fig. 2(a), the number of plausible hypotheses drops sharply as their length increases. This steep decline severely constrains our ability to apply structural complexity controls, as the available hypothesis space for longer, more complex structures becomes extremely limited. When attempting to simultaneously satisfy semantic content controls or structural complexity requirements, the intersection of valid hypotheses may become vanishingly small. (ii) Hypothesis Oversensitivity: As shown in Figure 2(b), when generating long logical hypotheses under control conditions, even small errors can result in significant differences in the conclusions. Previous work akgr used the Jaccard index as a reward in reinforcement learning, enforcing a perfect match between the hypothesis conclusion and the observation. This strict requirement may cause the model to deviate from the control conditions in order to maintain reward stability.
@@
To tackle these challenges, we propose a Controllable logcial Hypothesis Generation method (CtrlHGen) for abductive reasoning in knowledge graphs. To address the problem of hypothesis space collapse, we introduce a dataset augmentation strategy based on sub-logical decomposition. By leveraging the semantic similarity of simpler sub-logics derived from the decomposition of complex hypotheses, this approach enables the model to understand long logical structures, which are composed of these smaller components. The hypothesis generator is then trained using a combination of supervised fine-tuning and reinforcement learning. To address the problem of hypothesis oversensitivity, we refine the semantic reward function by incorporating the Dice coefficient to smooth out minor discrepancies between the hypothesis and the target. Additionally, we introduce a condition-adherence reward to encourage the generation of hypotheses that adhere to the control constraints during exploration. Our main contributions are as follows:
@@
  • To mitigate hypothesis oversensitivity, we refine the semantic reward function by incorporating Dice and Overlap coefficients to accommodate minor discrepancies between hypotheses and targets, while introducing a condition-adherence reward to ensure better compliance with control constraints, leading to more stable and accurate learning.
@@ 2 Related Work
Knowledge Graph Reasoning. Deductive reasoning focuses on answering complex logical queries by improving query and answer embeddings multihop ; query2box ; query2particles ; bai2023complex ; bai2023knowledge ; bai2024understanding . Inductive reasoning, often framed as rule mining, ranges from efficient symbolic methods like AMIE amie to embedding-based approaches such as RuLES rules and RLogic Rlogic , though traditional search-based techniques face scalability challenges. Abductive reasoning was introduced by AbductiveKGR akgr using Transformer-based hypothesis generation, with follow-up work ngdb highlighting its future potential.
Abductive Reasoning. In natural language inference, $\alpha$-NLI bhagavatula2020abductivecommonsensereasoning introduced abductive reasoning to commonsense reasoning, where plausible explanations are inferred from observations. Subsequent works proposed various techniques to enhance this capability qin2021futureunsupervisedbackpropbaseddecoding ; kadiķis2022embarrassinglysimpleperformanceprediction ; chan2023selfconsistentnarrativepromptsabductive , including extensions to uncommon scenarios focusing on rare but logical explanations zhao2024uncommonsensereasoningabductivereasoning . Unlike real-world data in commonsense reasoning, benchmarks like ProofWriter tafjord2021proofwritergeneratingimplicationsproofs evaluate formal abductive reasoning within semi-structured texts with explicit logical relationships. Recent studies have explored LLMs in more challenging open-world reasoning contexts zhong2023chatablabductivelearningnatural ; del2023truedetectivedeepabductive ; thagard2024chatgptmakeexplanatoryinferences and abstract reasoning tasks liu2024incompleteloopinstructioninference ; zheng2025logidynamicsunravelingdynamicslogical .
Meanwhile, in the neuro-symbolic domain, Abductive Learning (ABL) ABL attempts to integrate machine learning and logical reasoning in a balanced and mutually supportive manner. Recent research in this area, exemplified by systems such as ARLC ABL2 and ABL-Refl ABL3 , focuses on enhancing this integration by introducing novel techniques to improve context-awareness, error correction, generalization, and overall reasoning accuracy and efficiency.
@@ 3.1 Problem Definition
We define a knowledge graph as $G=(V,R)$, where $V$ is the set of entities and $R$ is the set of binary relations. A triple $(u,r,v)$ exists in $G$ if $r(u,v)=\text{true}$. Following the open-world assumption openworld , only the observed graph $G$ is available during training, with missing triples treated as unknown rather than false. The full graph $\bar{G}$ remains hidden, and $G\subseteq\bar{G}$.
The core concepts of abductive reasoning consist of observation and hypothesis. Here, an observation $O$ in knowledge graph $G$ is defined as a set of entities $O=\{o_{1},o_{2},\ldots,o_{n}\}$, where $o_{i}\in V,\forall i\in\{1,\ldots,n\}$ A logical hypothesis $H$ is defined as a query in the form of first-order logic on a knowledge graph $G$, including existential quantifiers($\text{And}(\wedge),Or(\vee),Not(\neg)$). The hypothesis can also be written in disjunctive normal form:
@@
The conclusion of the hypothesis $[H]_{G}$ on a graph $G$ is the set of the variable entities $V_{?}$ for which $H$ holds true on $G$. Specifically, it can be formulated as:
$[H]_{G}=\{V_{?}\in G|H(V_{?})=\mathrm{true}\}.$ (2)
@@
    The hypothesis $H$ is the most plausible explanation for the observation $O$ In other words, the conclusion $[H]_{G}$ closely matches the observation $O$.
@@ 3.2 Observation-Hypothesis Pairs Construction
Sampling. We randomly sample observation-hypothesis pairs from the knowledge graph by constructing hypotheses based on predefined logical patterns. Each logical pattern is assigned an equal number of hypotheses to ensure diversity, and the query results(conclusion) of hypotheses on the graph are taken as the corresponding observations. Finally, both hypotheses and observations are converted into input sequences suitable for the generative model.
@@
where $f(P_{\text{sub}}^{i},H)$ denotes the sub-hypothesis generated based on the sub-pattern $P_{\text{sub}}^{i}$ and the origin hypothesis $H$, and $[f(P_{\text{sub}}^{i},H)]_{G}$ computes the corresponding sub-observation by querying the knowledge graph to get the conclusion of the sub-hypothesis.
Because each sub-hypothesis is a subset of the original, they are closely related both structurally and semantically. This strong alignment enables the model to progressively learn complex logical patterns by building on simpler, related sub-patterns. We have reported more details and examples for the observation-hypothesis pairs construction process in Appendix A.
@@ 3.3 Supervised Training of Controllable Hypothesis Generation
  • Structural Constraint: We apply constraints based on the logic structure of the hypothesis. Specifically, we implement three types of structural control: (1) strictly enforcing a predefined logical pattern, where each logical pattern is represented in Lisp-like language with operator tokens following previous work in KG reasoning baisequential ; akgr . (2) constraining the number of entities involved, encoded using a special token [ne] that indicates hypotheses with exactly $n$ entities. Formally, $C\in\{[ne]\}$, where $n$ is an Integer. (3) constraining the number of relations used in the generated hypothesis, encoded using a token [nr], where [nr] denotes hypotheses containing exactly $n$ relations. Formally, $C\in\{[nr]\}$, where $n$ is an Integer.
@@ 3.4 Reinforcement Learning
Semantic Alignment: This reward assesses the semantic consistency between the generated hypothesis conclusion $[H]_{G}$ and the corresponding observation $O$. We adopt the Jaccard similarity coefficient as the primary reward due to its strict evaluation of set-level agreement. However, the high sensitivity of hypotheses can lead to sharp reward fluctuations in response to minor errors. To mitigate this, we integrate two supplementary metrics:the Dice similarity coefficient and the Overlap similarity coefficient, which provide smoother gradients and greater tolerance to slight mismatches. The final semantic reward $R_{\text{sem}}$ is computed as a weighted combination of these metrics, defined as:
@@
Since abductive reasoning often involves generating multiple plausible hypotheses rather than a single answer, it is important to ensure overall hypothesis quality. To achieve this, we use Group Relative Policy Optimization (GRPO) GRPO , which promotes consistent improvement across a set of sampled hypotheses per observation, instead of optimizing individual outputs. Specifically, GRPO updates the model $\pi_{\theta}$ by maximizing the expected reward over a group of hypotheses $\hat{H}={H_{1},\ldots,H_{k}}$ sampled from the same observation $O$ and control condition $C$. The objective is:
@@ 4.1 Experiment Settings
Dataset. We conduct experiments on three widely used knowledge graph datasets: DBpedia50 dbpedia , WN18RR wn18rr , and FB15k-237 fb15k . Following akgr , each dataset is split into training, validation, and test sets with an 8:1:1 ratio. Under the open-world assumption, we incrementally build $G_{\text{train}}$, $G_{\text{valid}}$, and $G_{\text{test}}$, where each graph includes all previous edges. Additional dataset details are provided in Appendix B.
Observation-Hypothesis Pair. Following prior KG reasoning work akgr , we adopt the 13 predefined logical patterns in Fig. 4 for hypothesis sampling. Each observation contains no more than 32 entities. To evaluate generalization, the validation and test sets include entities not seen during training, with the test set covering more unseen entities. For data augmentation, we decompose five complex logical patterns (up, 3in, pni, pin, inp). Detailed sampling statistics are provided in Appendix B.
@@
Evaluation Metrics. The quality of generated hypotheses is evaluated in terms of semantic similarity and condition adherence. For semantic similarity, we use Jaccard, Dice and Overlap index, with $G_{\text{test}}$ used to compute $[H]_{G_{\text{test}}}$ during testing. For condition adherence, we regard it as a binary classification problem and calculate Accuracy. In addition, Smatch score is also used to quantify the structural similarity corresponding to the generated hypothesis $H$ and the reference hypothesis $H_{\text{ref}}$. It can measure how similar the nodes, edges and their labels are by representing the hypothesis as a graph. It should be noted that Smatch is only a reference metric, as the generated hypotheses do not need to be the same as the reference hypotheses. More information about Smatch is reported in Appendix B.
Implementation Details. We adopt a GPT-2 gpt2 architecture for the hypothesis generation model and use the AdamW optimizer. All experiments are conducted on 4 Nvidia A6000 48GB GPUs. Additional hyperparameter settings and a discussion on why LLMs are not employed are provided in Appendix B  [^1] The code is available at https://github.com/HKUST-KnowComp/CtrlHGen ..

### v2

@@ Abstract
Abductive reasoning in knowledge graphs aims to generate plausible logical hypotheses from observed entities, with broad applications in areas such as clinical diagnosis and scientific discovery. However, due to a lack of controllability, a single observation may yield numerous plausible but redundant or irrelevant hypotheses on large-scale knowledge graphs. To address this limitation, we introduce the task of controllable hypothesis generation to improve the practical utility of abductive reasoning. This task faces two key challenges when controlling for generating long and complex logical hypotheses: hypothesis space collapse and hypothesis reward oversensitivity. To address these challenges, we propose CtrlHGen, a Controllable logcial Hypothesis Generation framework for abductive reasoning over knowledge graphs, trained in a two-stage paradigm including supervised learning and subsequent reinforcement learning. To mitigate hypothesis space collapse, we design a dataset augmentation strategy based on sub-logical decomposition, enabling the model to learn complex logical structures by leveraging semantic patterns in simpler components. To address hypothesis reward oversensitivity, we incorporate smoothed semantic rewards including Dice and Overlap scores, and introduce a condition-adherence reward to guide the generation toward user-specified control constraints. Extensive experiments on three benchmark datasets demonstrate that our model not only better adheres to control conditions but also achieves superior semantic similarity performance compared to baselines. Our code is available at https://github.com/HKUST-KnowComp/CtrlHGen.
@@ 1 Introduction
Abduction is widely recognized as one of the three major types of reasoning in philosophy (Douven, 2011). Specifically, abductive reasoning (Douven, 2011) is a form of logical inference that seeks the best or most plausible hypothesis to explain an observed phenomenon and it plays a vital role across various fields (Paul, 1993). For example, it serves as a critical tool for hypothesizing causal links between symptoms and underlying pathologies in clinical diagnosis (Pukancová and Homola, 2015; Martini, 2023). Similarly, abductive methods localize system faults by interpreting anomalous signal patterns in anomaly detection (Ramkumar et al., 2024; Ganesan et al., 2019). Its power also extends to scientific discovery (Engelschalt et al., 2023; Wackerly, 2021; Duede and Evans, 2021; Upmeier zu Belzen et al., 2021), including the deduction of unknown celestial bodies from gravitational perturbations in orbital trajectories (Smart, 1946).
@@
On the other hand, effective abductive reasoning requires high-quality, interconnected information. While large language models perform well in common-sense settings (Patil and Jadon, 2025), they often struggle in domains such as healthcare, business, or other scenarios involving sensitive data and strict privacy constraints. Knowledge graphs, whether general-purpose or domain-specific, provide a structured foundation that supports more reliable abductive reasoning. In knowledge graphs, abductive reasoning aims to generate complex logical hypotheses that explain observed entities, leveraging domain knowledge to improve inference precision and reliability. AbductiveKGR (Bai et al., 2024b) was the first to introduce this task, formulating it as logical query generation over structural knowledge and training models through a supervised–reinforcement learning framework.
However, knowledge graphs often contain millions of facts, which can lead to generate numerous plausible but irrelevant hypotheses from a single observation. For instance, even the relatively small DBpedia50 dataset(with only 24,624 entities and 351 relations), produces an average of 50 reasonable hypotheses per observation. In larger graphs, this number grows dramatically, underscoring the need to filter hypotheses according to user intent or interests for effective abductive reasoning. To address this challenge, we introduce controlling mechanisms into the hypothesis generation process, focusing on two critical aspects:
@@
Controlling structural complexity adjusts the level of granularity. We focus on complexity control to address varying information needs across different reasoning scenarios and align with users’ cognitive preferences for adjustable information density. In Fig. 1(b), for an observation composed of three NBA players, increasing the complexity of the hypothesis structure enables the model to capture richer shared experiences or achievements among them. By adjusting the structural complexity, users can flexibly decide how much information they want to include in the generated hypotheses. Unfortunately, prior work (Bai et al., 2024b) has largely overlooked controllable generation, resulting in hypotheses that are redundant or lack meaningful relevance.
@@
Motivated by these, we introduce the task of controllable abductive reasoning, aiming at controllable generation of hypothesis, which leads to better leverage the practical value of abductive reasoning in knowledge graphs. However, when implementing semantic and structural controls on complex long logical hypotheses, we face two critical challenges: (i) Hypothesis Space Collapse: As illustrated in Fig. 2(a), the number of plausible hypotheses drops sharply as their length increases. This sharp decline severely limits our ability to apply structural complexity control, as the model needs to ensure a strong understanding of complex logic in order to make correct candidate hypotheses. (ii) Hypothesis Reward Oversensitivity: The previous approach (Bai et al., 2024b) utilized the Jaccard score as a reward mechanism to enhance the model’s understanding of query semantics. However, as illustrated in Fig. 2(b), during the model’s exploration process, even a minor misstep may lead to a sharp drop in the Jaccard score, severely disrupting training stability and guiding the model toward incorrect directions.
@@
To tackle these challenges, we propose a Controllable logcial Hypothesis Generation method (CtrlHGen) for abductive reasoning in knowledge graphs. To address the problem of hypothesis space collapse, we introduce a dataset augmentation strategy based on sub-logical decomposition. By leveraging the semantic similarity of simpler sub-logics derived from the decomposition of complex hypotheses, this approach enables the model to understand long logical structures, which are composed of these smaller components. The hypothesis generator is then trained using a combination of supervised fine-tuning and reinforcement learning. To address the problem of hypothesis reward oversensitivity, we refine the semantic reward function by incorporating Dice and Overlap coefficient to smooth out minor discrepancies between the hypothesis and the target. Additionally, we introduce a condition-adherence reward to encourage the generation of hypotheses that adhere to the control constraints during exploration. Our main contributions are as follows:
@@
  • To mitigate hypothesis reward oversensitivity, we refine the semantic reward function by incorporating Dice and Overlap coefficients to accommodate minor discrepancies between hypotheses and targets, while introducing a condition-adherence reward to ensure better compliance with control constraints, leading to more stable and accurate learning.
@@ 2 Related Work
Knowledge Graph Reasoning. Deductive reasoning focuses on answering complex logical queries by improving query and answer embeddings (Zhang et al., 2021; Ren et al., 2020; Bai et al., 2022; 2023a; 2023b; 2024a). Inductive reasoning, often framed as rule mining, ranges from efficient symbolic methods like AMIE (Galárraga et al., 2013) to embedding-based approaches such as RuLES (Ho et al., 2018) and RLogic (Cheng et al., 2022), though traditional search-based techniques face scalability challenges. Abductive reasoning was introduced by AbductiveKGR (Bai et al., 2024b) using Transformer-based hypothesis generation, with follow-up work (Bai et al., 2025) highlighting its future potential.
Abductive Reasoning. In natural language inference, $\alpha$-NLI (Bhagavatula et al., 2020) introduced abductive reasoning to commonsense reasoning, where plausible explanations are inferred from observations. Subsequent works proposed various techniques to enhance this capability (Qin et al., 2021; kadiķis2022embarrassinglysimpleperformanceprediction; Chan et al., 2023), including extensions to uncommon scenarios focusing on rare but logical explanations (Zhao et al., 2024). Unlike real-world data in commonsense reasoning, benchmarks like ProofWriter (Tafjord et al., 2021) evaluate formal abductive reasoning within semi-structured texts with explicit logical relationships. Recent studies have explored LLMs in more challenging open-world reasoning contexts (Zhong et al., 2023; Del and Fishel, 2023; Thagard, 2024) and abstract reasoning tasks (Liu et al., 2024b; Zheng et al., 2025).
Meanwhile, in the neuro-symbolic domain, Abductive Learning (ABL) (Zhou, 2019) attempts to integrate machine learning and logical reasoning in a balanced and mutually supportive manner. Recent research in this area, exemplified by systems such as ARLC (Camposampiero et al., 2024) and ABL-Refl (Hu et al., 2025), focuses on enhancing this integration by introducing novel techniques to improve context-awareness, error correction, generalization, and overall reasoning accuracy and efficiency.
@@ 3.1 Problem Definition
We define a knowledge graph as $G=(V,R)$, where $V$ is the set of entities and $R$ is the set of binary relations. A triple $(u,r,v)$ exists in $G$ if $r(u,v)=\text{true}$. Following the open-world assumption (Drummond and Shearer, 2006), only the observed graph $G$ is available during training, with missing triples treated as unknown rather than false. The full graph $\bar{G}$ remains hidden, and $G\subseteq\bar{G}$.
The core concepts of abductive reasoning consist of observation and hypothesis. Here, an observation $O$ in knowledge graph $G$ is defined as a set of entities $O=\{o_{1},o_{2},\ldots,o_{n}\}$, where $o_{i}\in V,\forall i\in\{1,\ldots,n\}$ . A logical hypothesis $H$ is defined as a query in the form of first-order logic on a knowledge graph $G$, including existential quantifiers($\exists$), And($\wedge$), Or($\vee$), Not($\neg$). The hypothesis can also be written in disjunctive normal form:
@@
The conclusion of the hypothesis $H]_{G}$ on a graph $G$ is the set of the variable entities $V_{?}$ for which $H$ holds true on $G$. Specifically, it can be formulated as:
$H]_{G}=\{V_{?}\in G|H(V_{?})=\mathrm{true}\}.$ (2)
@@
    The hypothesis $H$ is the most plausible explanation for the observation $O$ . In other words, the conclusion $H]_{G}$ closely matches the observation $O$.
@@ 3.2 Observation-Hypothesis Pairs Construction
Sampling. We randomly sample observation-hypothesis pairs from the knowledge graph by constructing hypotheses based on predefined logical patterns. Each logical pattern is assigned an equal number of hypotheses to ensure diversity, and the conclusion of hypotheses on the graph are taken as the corresponding observations. Finally, both hypotheses and observations are converted into input sequences suitable for the generative model.
@@
where $f(P_{\text{sub}}^{i},H)$ denotes the sub-hypothesis generated based on the sub-pattern $P_{\text{sub}}^{i}$ and the origin hypothesis $H$, and $f(P_{\text{sub}}^{i},H)]_{G}$ computes the corresponding sub-observation by querying the knowledge graph to get the conclusion of the sub-hypothesis.
Because each sub-hypothesis is a subset of the original, they are closely related both structurally and semantically. This strong alignment enables the model to progressively learn complex logical patterns by building on simpler, related sub-patterns. We have reported more details in Appendix A.
@@ 3.3 Supervised Training of Controllable Hypothesis Generation
  • Structural Constraint: We apply constraints based on the logic structure of the hypothesis. Specifically, we implement three types of structural control: (1) strictly enforcing a predefined logical pattern, where each logical pattern is represented in Lisp-like language with operator tokens following previous work in KG reasoning (Bai et al., ; Bai et al., 2024b). (2) constraining the number of entities involved, encoded using a special token [ne] that indicates hypotheses with exactly $n$ entities. Formally, $C\in\{[ne]\}$, where $n$ is an Integer. (3) constraining the number of relations used in the generated hypothesis, encoded using a token [nr], where [nr] denotes hypotheses containing exactly $n$ relations. Formally, $C\in\{[nr]\}$, where $n$ is an Integer.
@@ 3.4 Reinforcement Learning
Semantic Alignment: This reward assesses the semantic consistency between the generated hypothesis conclusion $H]_{G}$ and the corresponding observation $O$. We adopt the Jaccard similarity coefficient as the primary reward due to its strict evaluation of set-level agreement. However, the high sensitivity of hypotheses can lead to sharp reward fluctuations in response to minor errors. To mitigate this, we integrate two supplementary metrics:the Dice similarity coefficient and the Overlap similarity coefficient, which provide smoother gradients and greater tolerance to slight mismatches. The final semantic reward $R_{\text{sem}}$ is computed as a weighted combination of these metrics, defined as:
@@
Since abductive reasoning often involves generating multiple plausible hypotheses rather than a single answer, it is important to ensure overall hypothesis quality. To achieve this, we use Group Relative Policy Optimization (GRPO) (Shao et al., 2024), which promotes consistent improvement across a set of sampled hypotheses per observation, instead of optimizing individual outputs. Specifically, GRPO updates the model $\pi_{\theta}$ by maximizing the expected reward over a group of hypotheses $\hat{H}={H_{1},\ldots,H_{k}}$ sampled from the same observation $O$ and control condition $C$. The objective is:
@@ 4.1 Experiment Settings
Dataset. We conduct experiments on three widely used knowledge graph datasets: DBpedia50 (Auer et al., 2007), WN18RR (Bordes et al., 2013), and FB15k-237 (Toutanova and Chen, 2015). Following (Bai et al., 2024b), each dataset is split into training, validation, and test sets with an 8:1:1 ratio. Under the open-world assumption, we incrementally build $G_{\text{train}}$, $G_{\text{valid}}$, and $G_{\text{test}}$, where each graph includes all previous edges.
Observation-Hypothesis Pair. Following prior KG reasoning work (Ren et al., 2020), we adopt the 13 predefined logical patterns in Fig. 4 for hypothesis sampling. Each observation contains no more than 32 entities. To evaluate generalization, the validation and test sets include entities not seen during training, with the test set covering more unseen entities. For sub-logic decomposition, we chose five complex logical patterns (up, 3in, pni, pin, inp) to break down.
@@
Evaluation Metrics. The quality of generated hypotheses is evaluated in terms of semantic similarity and condition adherence. For semantic similarity, we use Jaccard, Dice and Overlap score, with $G_{\text{test}}$ used to compute $H]_{G_{\text{test}}}$ during testing. For condition adherence, we regard it as a binary classification problem and calculate Accuracy. In addition, Smatch score (Cai and Knight, 2013) is also used to quantify the structural similarity corresponding to the generated hypothesis $H$ and the reference hypothesis $H_{\text{ref}}$. It can measure how similar the nodes, edges and their labels are by representing the hypothesis as a graph. It is an evaluation metric for Abstract Meaning Representation (AMR) graphs, which are directed acyclic graphs with two node types (variable and concept) and three edge types (instance, attribute, and relation). Given a predicted graph $G_{p}$ and a gold graph $G_{g}$, Smatch($G_{p},G_{g}$) is computed by finding an approximately optimal mapping between the variable nodes of the two graphs and matching their edges. Following the settings of  Bai et al. (2024b), we transform the hypothesis graph $G(H)$ into an AMR graph $GA(H)$ by adding virtual nodes and instance edges, and then calculate Smatch. In short, Smatch is used to measure the degree of similarity between the generated hypothesis and the ground truth in the test set. It should be noted that Smatch is only a reference metric, as the generated hypotheses do not need to be the same as the reference hypotheses.
Implementation Details. We adopt a 12 layers decoder-only Transformer architecture (Radford et al., 2019; Vaswani et al., 2017) for the hypothesis generation model and use the AdamW optimizer. All experiments are conducted on 4 Nvidia A6000 48GB GPUs. Additional hyperparameter settings and other experiment details are reported in Appendix B.
@@
Table 2: Average scores on FB15k237 datasets under five conditions
Model Jaccard Dice Overlap Accuracy Smatch
GPT-4o + 2-hop subgraph 2.4±0.10 3.1±0.13 5.3±0.20 77.5±0.31 37.9±0.27
Kimi K2 + 2-hop subgraph 3.1±0.11 4.7±0.17 8.5±0.24 71.6±0.34 42.4±0.22
Grok-3 + 2-hop subgraph 2.5±0.09 3.7±0.12 6.9±0.21 75.6±0.38 43.5±0.21
DeepSeek-V3 + 2-hop subgraph 2.1±0.09 2.8±0.11 6.3±0.26 73.9±0.33 41.8±0.27
DeepSeek-V3 + RAG 5.3±0.15 6.7±0.17 10.4±0.46 76.6±0.35 41.8±0.27
GPT5(Thinking) + 2-hop subgraph 18.7±0.32 21.9±0.35 37.3±0.46 92.8±0.28 32.9±0.27
CtrlHGen 64.3±0.33 71.9±0.31 83.0±0.29 96.6±0.84 73.3±0.16